Raymond Mining Corporation has 8.8 million shares of common stock outstanding, 320,000 shares of 4% $100 par value preferred stock outstanding, and 149,000 7.50% semiannual bonds outstanding, par value $1,000 each. The common stock currently sells for $36 per share and has a beta of 1.40, the preferred stock currently sells for $92 per share, and the bonds have 10 years to maturity and sell for 117% of par. The market risk premium is 7.6%, T-bills are yielding 5%, and Raymond Mining’s tax is 38%. The YTM before taxes is 5.289%.

What is the WACC? (Enter your answers in whole dollars. Omit % sign in your response.)

q25
Answer #1

Value Debt $174,330,000.00 Common stock $316,800,000.00 Preferred $ 29,440,000.00 $520,570,000.00 Weight Cost Weight* cost 33

Calc:

Debt Common stock Preferred Value =149000* (1000*117%) =8800000*36 =320000*92 =SUM(C11:C13) Weight =C11/C14 =C12/C14 =C13/C14